OpenStudy (anonymous):

PLZ HELP ME!!! Part A: Max rented a motorbike at $465 for 5 days. If he rents the same motorbike for a week, he has to pay a total rent $625. Write an equation in the standard form to represent form to represent the total rent (y) that Max has to pay for renting the motorbike for x days. Part B: Write the equation obtained in part A using function notation. Part C: Describe the steps to the graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals.

OpenStudy (anonymous):

PLZ Z HELP I HAVE BEEN STUCK FOR THE PAST HOUR

OpenStudy (anonymous):

FUNCTION NOTATION IS f(x) = mx + b

OpenStudy (anonymous):

here this is kinda like that problem (03.06 MC) Part A: Michael rented a car for 2 days at $225.00 per day. If he rents the same car for 5 days, he has to pay a total of $480. Write an equation in the standard form that represent the total rent (y) that Michael has to pay for renting the car for x days. (4 points) Part B: Write the equation obtained in Part A using function notation. (2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points) (10 points) Part A: Taking that he paid 480 for 5 days .. states he pays 96 a day, y = 96x for the "x" days of renting a car. Goes with theory that if you keep the car longer your day rate will be cheaper Part B: f(x) = 96x Part C: you plot two points and make a line between them

OpenStudy (anonymous):

answer is the second a,b and c

OpenStudy (anonymous):

Thank you SOOO MUCH that was a big help

OpenStudy (anonymous):

Do u need anything else?

OpenStudy (anonymous):

yes see we helped eachother i have two questions one sec

OpenStudy (anonymous):

12. (03.06 MC) Which equation can you write for a line that will pass through the points (3, 6) and (5, 18) in function notation. (5 points) f(x) = 2x + 12 f(x) = 6x − 12 y = 2x + 12 y = 6x − 12 Score: 0 of 5 13. (03.07 LC) Choose the equation of the horizontal line that passes through the point (−5, 2). (5 points) y = −5 y = 2 x = −5 x = 2 Score: 0 of 5